# A Cost-Efficient FPGA Implementation of Tiny Transformer Model using Neural ODE

### v2

@@ 6.2.6 Execution time
In this section, the performance of the FPGA implementation is compared with its software counterpart running on ARM Cortex-A53 CPU. The software counterpart is implemented with PyTorch. The 8-bit and 4-bit quantizations are applied to both the implementations, respectively. Table 7 presents the mean, maximum, and standard deviation of the execution time obtained by 100 runs. Please note that, in the FPGA implementation, only the pre- and post-processing is executed on the PS part while the feature extraction part is executed on the PL part. The total execution times (PS + PL) of the 8-bit and 4-bit FPGA implementations are 162.1ms and 161.8ms; thus, they achieve 9.85$\times$ and 9.83$\times$ speedups, respectively. If we look at only the feature extraction part, they achieve 34.01$\times$ and 33.96$\times$ speedups, respectively.
Table 7 also presents GOPS (giga operations per second) values based on the mean execution times. The peak performance is 41.2 GOPS when 206 DSP slices are fully-utilized for multiply and add operations at 200MHz, while the measured GOPS value of the PL part is 4.33 GOPS. The proposed accelerator is implemented as a stream of separated blocks, such as ODEBlock1, DSBlock1, ODEBlock2, DSBlock2, and MHSABlock, as shown in Fig. 6. Although our current accelerator design is modular, meaning that these blocks can be added or removed depending on the model, it would be possible to bring the measured GOPS value closer to the peak GOPS value by reducing the number of DSP slices by reusing the same DSP slices for all the blocks. Such architectural exploration and optimization are our future work.
@@ 6.2.7 Energy efficiency
7 Summary
@@
Transformers, featuring the highly expressive MHSA mechanism, have demonstrated remarkable accuracy in the field of image processing among recent AI models. This paper focuses on a hybrid model that combines Transformer and CNN architectures. Considering the resource constraints of edge devices, we employed Neural ODE, an approximation technique inspired by ResNet, to design a tiny CNN-Transformer hybrid model with 95.1% less parameters than ResNet50. The parameter size of the proposed model is the smallest among the state-of-the-art lightweight models while it offers a comparable accuracy. We proposed the resource-efficient FPGA design of the tiny Transformer model tailored for edge devices. We employed a recently-proposed LUT-based quantization technique and conducted the ablation studies to investigate the impact of quantization on different parts of the proposed model. The entire feature extraction part is implemented on a modest-sized FPGA to accelerate inference speed. Compared to an embedded CPU, the FPGA implementation achieved a 34.01$\times$ faster runtime without pre- and post-processing and a 9.85$\times$ faster runtime for the overall inference, resulting in a 7.10$\times$ higher energy efficiency. Thanks to the small parameter size and quantization, the parameters fit within the limited on-chip memory, eliminating unnecessary data transfer overhead. Using high-end FPGAs would allow more design choices and further optimizations, potentially leading to greater improvements on the speed and accuracy. While basic forms of the attention mechanism and Neural ODE are employed in this paper, it is possible to further refine the proposed model in terms of computational efficiency and accuracy, as discussed in Sec 2. In this paper, the proposed model was implemented as a sequence of multiple blocks in the PL part. As an alternative approach, it can be implemented as a single configurable block. Although the single block implementation is more complex and challenging, it would be beneficial from a resource efficiency viewpoint, enabling a performance improvement in cooperation with more aggressive parallelization. A demonstration video of the proposed FPGA implementation on ZCU104 board is available on YouTube  [^2] https://youtu.be/jldJGnSYfVQ?si=Ss0PWq2i17hB2Q6S.

### v3

@@ 6.2.6 Execution time
In this section, the performance of the FPGA implementation is compared with its software counterpart running on ARM Cortex-A53 CPU. The software counterpart is implemented with PyTorch. The 8-bit and 4-bit quantizations are applied to both implementations. Table 8 presents a distribution of the execution time obtained by 100 runs. The results are visualized in Fig. 13. Please note that, in the FPGA implementation, only the pre- and post-processing part is executed on the PS part while the feature extraction part is handled in the PL part. The total execution times (PS + PL) of the 8-bit and 4-bit FPGA implementations are 162.1ms and 161.8ms; thus, they achieve 9.85$\times$ and 9.83$\times$ speedups, respectively. If we only consider the feature extraction part, they achieve 34.01$\times$ and 33.96$\times$ speedups, respectively.
Table 8 also presents GOPS (giga operations per second) values based on the mean execution times. The results are visualized in Fig. 14. The peak performance is 41.2 GOPS when 206 DSP slices are fully-utilized for multiply and add operations at 200MHz, while the measured GOPS value of the PL part is 4.33 GOPS. The proposed accelerator is implemented as a stream of diverse blocks, such as ODEBlock1, DSBlock1, ODEBlock2, DSBlock2, and MHSABlock, in contrast to classic ResNets that have very uniform structure. Considering this diverse structure, our current accelerator is implemented as a modular manner. In this case, it would be possible to bring the measured GOPS value closer to the peak GOPS value by reducing the number of DSP slices by reusing the same DSP slices for all the blocks. Also, it would be possible to further reduce the number of DSP slices for the 4-bit implementation by adopting an aggressive DSP packing technique. Such architectural exploration and optimization are our future work.
@@ 6.2.7 Energy efficiency
Figure 13: Execution time (ms) of CPU and FPGA implementations
Figure 14: GOPS of CPU and FPGA implementations
6.2.8 Performance comparison against existing work
In [13], a hybrid model that combines MHSA and Neural ODE based CNN is implemented on ZCU104, though only the MHSA part is accelerated by the PL part. The performance gain of the MHSA part is only 2.63$\times$ compared to the software counterpart [13]. In [39], a Neural ODE based CNN is implemented on the same platform. The performance gain without the pre- and post-processing is 27.9$\times$ [39] and its fps (frames per second) is 16.6. In this work, on the other hand, the performance gain of the MHSA and CNN parts without the pre- and post-processing is 34.01$\times$ and its fps is 23.0, resulting in a higher performance compared to these works on the same FPGA platform. Although our targets are resource-limited FPGA devices, ViT variants are accelerated by Xilinx Alveo U200 platform [33] and their ME-ViT achieves 11.99 to 23.98 fps for larger images.
6.2.9 Practical implications
Reducing the parameter size while preserving the accuracy is challenging while it is usually required in resource-limited edge computing. Below are our practical implications toward this challenge which were obtained through the experiments:
  1. 1.
    Our ablation studies show that the DS+Block3 model where the 8-bit quantization is applied to DSBlocks and MHSABlock strikes a good balance between the accuracy and parameter size.
  2. 2.
    Applying DSC to DSBlocks can further reduce the parameter size with small accuracy loss, showing a favorable trade-off between them.
  3. 3.
    Our quantized models are first initialized with parameters of the baseline floating-point model and then fine-tuned with a smaller learning rate. The LLT-based quantization method can be implemented on the FPGA with small overheads while mitigating the accuracy loss.
7 Conclusions
7.1 Summary
Transformers, featuring the highly expressive MHSA mechanism, have demonstrated remarkable accuracy in the field of image processing among recent AI models. This paper focuses on a hybrid model that combines Transformer and CNN architectures. Below are our key contributions:
  1. 1.
    Considering the resource constraints of edge devices, we employed Neural ODE, an approximation technique inspired by ResNet, to design a tiny CNN-Transformer hybrid model with 95.1% less parameters than ResNet50. The parameter size of the proposed model is the smallest among the state-of-the-art lightweight models while it offers a comparable accuracy.
  2. 2.
    We proposed the resource-efficient FPGA design of the tiny Transformer model tailored for edge devices. We employed a recently-proposed LUT-based quantization technique and conducted the ablation studies to investigate the impact of quantization on different parts of the proposed model.
  3. 3.
    The entire feature extraction part is implemented on a modest-sized FPGA to accelerate inference speed. As a result, the FPGA implementation achieved a 34.01$\times$ faster runtime without pre- and post-processing and a 9.85$\times$ faster runtime for the overall inference compared to an embedded CPU, resulting in a 7.10$\times$ higher energy efficiency.
A demonstration video of the proposed FPGA implementation on ZCU104 board is available on YouTube  [^4] https://youtu.be/jldJGnSYfVQ?si=Ss0PWq2i17hB2Q6S.
7.2 Limitations and future direction
@@
The objective of this work is to demonstrate our proposed lightweight hybrid model combining Neural ODE and MHSA on resource-limited FPGA devices for edge computing. Thanks to the adoption of Neural ODE and quantization, the parameter size is significantly reduced and thus the proposed model fits within the limited on-chip memory, eliminating unnecessary data transfer overhead. On the other hand, as mentioned in Sec. 6.2.8, the performance of the proposed IP core is not always higher than those for larger FPGA devices [33]. We believe that using larger FPGAs would allow more design choices and optimizations, potentially leading to greater improvements on the speed and accuracy. While basic forms of the attention mechanism and Neural ODE are employed in this paper, it is possible to further refine the proposed model in terms of computational efficiency and accuracy, as discussed in Sec. 2. In addition, in this paper the proposed model was implemented as a sequence of multiple blocks in the PL part. As an alternative approach, it can be implemented as a single configurable block. Although the single block implementation is more complex and challenging, it would be beneficial from a resource efficiency viewpoint, enabling a performance improvement in cooperation with more aggressive parallelization.